OpenStudy (anonymous):

5a^2+23a+12 factor this, show work, i dont know how to do this

OpenStudy (shamim):

23=20+3 right?

OpenStudy (triciaal):

do you know how to complete the square or do you know the quadratic formula?

OpenStudy (shamim):

20*3=60

OpenStudy (anonymous):

quadratic formula

OpenStudy (shamim):

5a^2+20a+3a+12 right?

OpenStudy (anonymous):

i guess, why did u do the 20a+3a?

OpenStudy (triciaal):

@shamin I think you should start at the beginning to show why use 20 and 3

OpenStudy (triciaal):

see

OpenStudy (triciaal):

@shamim

OpenStudy (shamim):

20a+3a=23a right?

OpenStudy (anonymous):

i got -16.2 and -29.5 from using the quadratic formula

OpenStudy (shamim):

20a*3a=60a^2

OpenStudy (shamim):

nd 5a^2*12=60a^2

OpenStudy (triciaal):

@splatter6 @shamim is going through how to factor follow along for a minute @shamim please go through all the steps

OpenStudy (anonymous):

ok

OpenStudy (shamim):

understood?

OpenStudy (anonymous):

no, i dont understand y u have a 60

OpenStudy (shamim):

60a^2=60a^2 right?

OpenStudy (triciaal):

@shamim please wait @splatter6

OpenStudy (shamim):

need to show u a drawing

OpenStudy (shamim):

bt i m now frm an android phone. so not possible to draw

OpenStudy (shamim):

sorry

OpenStudy (anonymous):

it's fine, i'll just ask my teacher

OpenStudy (shamim):

which grade u r

OpenStudy (anonymous):

im 11th

OpenStudy (shamim):

r u frm usa?

OpenStudy (anonymous):

yeah

OpenStudy (shamim):

ok

OpenStudy (anonymous):

y

OpenStudy (shamim):

do u know middle term for factorisation?

OpenStudy (anonymous):

no

OpenStudy (triciaal):

when you have a quadratic in the form ax^2 + bx + c to factor find the product a*c and then the factors of that with the combination such that the sum will be the coefficient of x namely b for this problem a = 5 b = 23 an d c = 12 so 5*12 = 60 what are the factors of 60 that when you add them the sum is 23 factors of 60 eg 6 and 10 , 15 and 4, 5 and 12 and the one used 20 and 3 now you have 5x^2 + 20x + 3x + 12 = 0 5x( x + 4) + 3( x + 4) = 0 @shamim will continue

OpenStudy (shamim):

i used it

OpenStudy (anonymous):

ok